OpenStudy (anonymous):

1 + sec^2x sin^2x = sec^2x

OpenStudy (loser66):

prove?

OpenStudy (anonymous):

yeah

OpenStudy (loser66):

\[sec^2 x= \frac{1}{cos^2 x}\] so plug there to get \[1+\frac{sin^2x}{cos^2 x}= 1 + tan^2 x = sec^2 x\]that 's it

OpenStudy (anonymous):

thanks!

OpenStudy (loser66):

ok,